OpenStudy (anonymous):

negative 1 over 2 x + 2 = −x + 7

OpenStudy (anonymous):

is this the equation? \[-\frac{ 1 }{ 2x+2 }= -x+7\]

OpenStudy (anonymous):

-1 = (2x + 2)(-x + 7) -1 = -2x^2 +14x - 2x + 14 2x^2 - 12x - 15x = 0 Quadratic formula from there would be easiest...